OpenStudy (anonymous):

3.04]What is the Least Common Denominator (LCD) of eleven-fourteenths and sixteen twenty-firsts? 7 fourteen thirty-two forty-two

OpenStudy (anonymous):

42

OpenStudy (anonymous):

thanks meadal an fan here u go @juliannescriv

OpenStudy (anonymous):

find the lowest common multiple of 14 and 21. 14 = 7 * 2 21 = 7 * 3 so least common multiple = 7 * 2 * 3